# Radio over Fiber with Cascaded Structure: Algorithm for Uplink Positioning

### v1

@@ Abstract
Recent advancements in polymer microwave fiber (PMF) technology have created significant opportunities for robust, low-cost, and high-speed sub-terahertz (THz) radio-over-fiber communications. Recognizing these potential benefits, this paper explores a novel radio-over-fiber (RoF) structure that interconnects multiple radio units (RUs) in cascade via fiber, envisioning its application in indoor scenarios. This structure creates a number of research opportunities when considering cascaded distortion effects introduced by non-linear power amplifiers (PAs) at the RUs and the propagation channel over the fiber.
We propose maximum-likelihood and non-linear least-squares algorithms to estimate the propagation distance along the RoF and the time-of-arrival between the RoF and the user equipment. For the case of linear PAs, we derive the Cramér-Rao lower bound to benchmark the performance of the estimators. Finally, we investigate the use of the system for uplink positioning. Our simulation results demonstrate that the proposed estimators perform satisfactorily even with the cascaded effects of non-linear PAs, and that the deployment of this RoF structure can enable new cost-effective opportunities for high-resolution positioning in indoor scenarios. In the numerical evaluation, we also use measured PMF characteristics for high-density polyethylene fibers.
@@ I Introduction
Next-generation wireless communication systems are expected to support ubiquitous wireless connectivity. To achieve this goal, one crucial limitation is the spectrum resources. Therefore, the opportunities to use the abundance of the spectrum at the sub-terahertz (sub-THz) frequency band have been identified as a promising enabler for both ultra-high speed communications [2, 3] and high-resolution positioning [4]. This is largely so because sub-THz wireless systems can make use of a large bandwidth, allowing high data rates (up to several terabits per second) and high spatial resolution.
Although there has been abundant research on the performance of sub-THz wireless communication systems [5, 6, 7], their implementation is still challenging [8]. Considering the cost of sub-THz radio hardware and the limited coverage [9], a low-cost and densely deployable solution is desirable. Recent developments in polymer microwave fiber (PMF) have motivated the development of radio-over-fiber (RoF) communications, which offers great opportunities for low-cost implementations of sub-THz wireless systems [10, 11, 12]. RoF systems can be implemented in different ways. Herein, we consider a system with a cascaded structure, where multiple radio units (RUs) are interconnected via low-cost PMFs and finally connected to a central unit (CU), where all signal processing is performed. Fig. 1 illustrates the cascade RoF system under consideration (the different variables will be introduced later). In this system, the signal propagates over a short-range sub-THz wireless channel and over multiple dispersive fibers within the RoF system. The salient feature of this system is that the propagation distance over the sub-THz wireless channel is only a small fraction of the distance between the CU and the user equipment (UE). Consequently, the overall propagation attenuation is significantly mitigated compared to that of a pure wireless sub-THz system of the same coverage. According to measurements reported in [13], PMFs attenuate sub-THz signals more than 2 dB per meter in the sub-THz spectrum. Hence, RUs have to be densely deployed to maintain the signal strength. RUs can either serve as access points or boosters (that amplify and forward the signals); thus the closest RU to the UE acts as access point and all other RUs between the entry/exit one and the CU act as boosters to compensate for the attenuation over the fiber. The RUs are equipped with antennas, a power amplifier (PA), and other radio frequency (RF) components. Since there are not digital signal processing units at the RUs, the cost of the system is significantly less than the dense deployment of radio transceivers.
@@
For energy efficiency (EE) purposes, the PAs are allowed to operate in the non-linear regime [14]. Therefore, at each stage of the RoF system, the signals are distorted by a non-linear (PA) and a segment of dispersive PMF, with the latter being considered a linear time-invariant system. After multiple stages, the signal will be subjected a cumulative distortion, which we exploit to estimate the propagation distance over the RoF. Thus, in combination with time-of-arrival (TOA) estimation over the wireless channel, the proposed system enables high-resolution positioning of UEs.
@@
Contributions: The specific contributions of this paper are (i) an analysis of the cascaded structure with non-linear PAs and dispersive PMFs, and an analysis of the uplink (UL) signal; (ii) a maximum-likelihood (ML) framework and a non-linear least-squares (NLS) framework for the propagation distance estimation; (iii) a thorough evaluation on the performance of the estimators with PAs working in the linear regime via an analysis of Cramér-Rao lower bound (CRLB); (iv) numerical results (using measured PMF dispersion characteristics and different values of non-linear factors) demonstrate that the proposed estimators perform well on the propagation distance estimation even in the presence of the effects of cascaded non-linear PAs; and (v) we demonstrate the usability of the proposed system for high-resolution positioning in indoor scenarios. To the best of our knowledge, this is the first study of a RoF system with segments of dispersive PMFs and with non-linear amplifiers connected in cascade. The proposed system and analysis have the potential to open up several novel research directions and opportunities for positioning and localization systems.
The rest of this paper is organized as follows. Section II presents related works. Section III discusses the system overview and the structure of every component within a RoF. Section IV formulates the problem and the corresponding signal models. Section V evaluates the ML estimator in the linear regime by CRLB. Section VI shows the applied channel measurement data and numerical results. Section VII applies the proposed algorithm in an indoor positioning use case. Section VIII concludes the paper.
@@ II Related work
It is worth noting that this study considers a RoF system formed with only RF components and fibers. This implementation is novel and has not been investigated before, and it should not be confused with other RoF approaches that we introduce in the following.
Different from our proposal, a traditional RoF system considers the merging of radio frequency and optical fiber technologies, whose advantages are introduced in [15]. It is broadly acknowledged that RoF systems present low attenuation and high cost-efficiency due to the use of low-cost fibers whose development is reported in [10, 11, 12]. Moreover, since the signals in RoF systems are well confined within the fiber, these systems are not restricted in bandwidth. The work in [16] experimentally demonstrates the RoF systems in the THz frequency band, which implements a fiber-wireless seamless network in an indoor scenario. The reported results demonstrate good performance in terms of carrier-to-noise ratio. However, this type of RoF system often suffers from severe non-linearity in optical components, such as lasers [17], and materials[18], which represent the main challenge for RoF communications.
It is also worth differentiating the proposed system from radio stripes (RS), which is a commonly envisioned implementation of distributed massive MIMO (D-MMIMO). D-MMIMO distributes the antennas in a large area to provide ubiquitous service for every UE. The work in [19] provides a comprehensive overview of D-MMIMO in 6G. Although the D-MMIMO offers consistently high-quality service to all UEs, accurate synchronization between APs and the CU is challenging. On the other hand, RSs connect APs sequentially in one cable/fiber to the CU [20], which achieves simplified synchronization between APs via the shared bus. The authors in [21] investigate the high-accuracy positioning capability with RSs at the sub-6 GHz band. However, previous works about RS consider that every AP is equipped with full signal processing abilities, which are considerably expensive at THz frequencies. However, the unique structure considered in this paper allows dense and flexible deployment because the CU allocates all the resources and processes the signals. The studied system distinguishes itself from the distributed massive MIMO in the literature due to the absence of multiple antenna signal processing.
@@ III System Overview
Consider the system illustrated in Fig. 2, where multiple RoFs are deployed in parallel to serve UEs in an indoor scenario. Each RoF consists of $U$ RUs and one CU at the end. One central coordination unit (CCU) connects the RoFs with the servers and coordinates different RoFs. All RUs are installed in the ceiling at the same height. The RUs are uniformly distributed, meaning the distance between two consecutive RUs in one RoF is fixed as $\gamma$ and the distance between any two RoFs is also fixed as $\gamma$. The location of the UE is denoted by $(P_{x},P_{y},P_{z})$ and the location of the $r$th RU in the $m$th RoF is denoted by $(r\gamma,m\gamma,0)$. Therefore, the distance between the UE and the RUm,r is calculated as
$d_{m,r}=\sqrt{\lvert P_{x}-r\gamma\rvert^{2}+\lvert P_{y}-m\gamma\rvert^{2}+P_{z}^{2}}.$ (1)
The configurations of RU, CU, and PMF will be discussed below. Other components are all considered to be linear filters having the same effect as the PMF.
Figure 3: Forward or receive configuration of the RU.
Figure 4: Characteritics of the PA with different values of the non-linearity factor.
Figure 5: Central unit.
@@ III-A Polymer microwave fiber
For a unit-length fiber segment, the channel impulse response with $L$ taps and amplitudes ($\beta_{0},\beta_{1},\cdots,\beta_{L-1}]^{\mathrm{T}}$) [^1] This model is valid in a non-linear system if the signals are oversampled properly. can be expressed as
@@
where $n$ is the time index. Then, the impulse response of a fiber with $r$ unit length can be expressed as
@@
where all factors in (3) are modeled as in (2). In the frequency domain, we consider the frequency response for a set of $K$ discrete frequencies, $\boldsymbol{f}=[f_{0},\cdots,f_{K-1}]^{\mathrm{T}}$. Thus, the frequency domain representatioin of (3) is given by the multiplication of frequency responses. Let $\boldsymbol{H}$$H_{0},\cdots,H_{K-1}]^{\mathrm{T}}$ represent the frequency response of a unit-length fiber. Then, the frequency response of a unit-length fiber at the $k$th frequency $H_{k}$ is the Fourier transform of its impulse response given by
@@
III-B Radio Unit
RUs are responsible for capturing incoming sub-THz signals, amplifying them, and forwarding them to the next component in the same stripe. Each RU consists of PMF couplers, PAs, antennas, and a switch as illustrated in Fig. 3. For the operation of the RoF system, we consider that each RU is equipped with a switch that selectively receives signals, either from the PMF fibers or the sub-THz antennas. When the RU detects signals from the PMFs, the switch disconnects the antenna input, allowing only the PMF signals to pass through. Thus, when the RU is configured to receive only PMF signals, it operated purely as a booster. Consequently, each UE is served exclusively by the nearest RU.
To understand the impact of non-linearities, we examine both regimes of operation of PAs, linear and non-linear.. The effects of non-linear PAs are characterized as a third-order polynomial with a factor $\lambda$ [22, 23]. Therefore, relationship between the input $x_{n}$ and output $y_{n}$ of a PA is given by
@@ III-B Radio Unit
where $G$ is the amplification coefficient. Fig. 4 illustrates the characteristics of PA with different values of non-linear factors, where we can observe the non-linearity arises with a signal magnitude of $\sim 0.4$ volts.
It should be noteded that every RU is incapable of processing signals because of the absence of a digital signal processing unit. From the perspective of expenditure, this design significantly lowers the price of each RU, since a digital processing unit is expensive in high frequencies. Conversely, this design disables the RU to mitigate signal impairments introduced by PAs and PMFs. As a consequence, this results in the accumulation of signal distortion along the propagation in a RoF system.
@@ III-C Central unit
The CU is in charge of allocating resources, controlling the RUs, and processing data. Each CU comprises PAs, mixers, couplers, ADC/DAC, phase-locked loops (PLL), and a digital signal processing unit (see Fig. 5). The CU is capable to operate in full-duplex mode to communicate with one UE simultaneously.
@@ IV Problem formulation and signal model
Consider the indoor scenario illustrated in Fig. 2, where the UE transmits data to multiple RoFs through the UL communication, where all RUs are active. Following the discussion in Section III, this paper considers two RU configurations with PAs working in both linear and non-linear regimes. In this paper, we assume all PAs have identical characteristics and work at the same operating point to simplify the analysis. Herein, the signals would experience the same distortion at every stage.

### v2

@@ Abstract
Recent advancements in polymer microwave fiber (PMF) technology have created significant opportunities for robust, low-cost, and high-speed sub-terahertz (THz) radio-over-fiber communications. Recognizing these potential benefits, this paper explores a novel radio-over-fiber (RoF) structure that interconnects multiple radio units (RUs), booster units (BUs), and a central unit (CU) in cascade via fiber, envisioning its application in indoor scenarios. This structure creates a number of research opportunities when considering cascaded distortion effects introduced by non-linear power amplifiers (PAs) and the propagation channel over the fiber.
We propose maximum-likelihood and non-linear least-squares algorithms to estimate the entry RU and the time-of-arrival between the RoF and the user equipment, where estimating the entry RU is equivalent to estimating the propagation distance along the RoF stripe. For the case of linear PAs, we derive the Cramér-Rao lower bound to benchmark the performance of the estimators. Finally, we investigate the use of the system for uplink positioning. Our simulation results demonstrate that the proposed estimators perform satisfactorily even with the cascaded effects of non-linear PAs, and that the deployment of this RoF structure can enable new cost-effective opportunities for high-resolution positioning in indoor scenarios. In the numerical evaluation, we also use measured PMF characteristics for high-density polyethylene fibers.
@@ I Introduction
NEXT-generation wireless communication systems are expected to support ubiquitous wireless connectivity. To achieve this goal, one crucial limitation is the spectrum resources. Although sixth-generation (6G) communication systems are expected to adopt frequency bands that strike a balance between capacity and coverage, particularly within the new mid-band spectrum (7.125 to 24.25 GHz) [bazzi2025upper], the sub-terahertz (sub-THz) range (90 to 300 GHz) is anticipated to enable high-precision sensing and short-range positioning applications [THzsurvey, cai6G, 6Gwhitepaper]. Consequently, research on sub-THz technologies is expected to continue advancing [shafi2025industrial]. In this work, we present a novel system operating in the sub-THz band to support ultra-high data rates, low latency, and enhanced positioning accuracy. This is largely due to the wide bandwidth available in the sub-THz spectrum, which enables data transmission rates of several terabits per second and provides high spatial resolution.
Although there has been abundant research on the performance of sub-THz wireless communication systems [THzdata, THzXR, SubTHzPosition], their implementation is still challenging [THzimplementationchallenge]. Considering the cost of sub-THz radio hardware and the limited coverage [bjornson2019massive], a low-cost and densely deployable solution is desirable. Recent developments in polymer microwave fiber (PMF) have motivated the development of radio-over-fiber (RoF) communications, which offer great opportunities for low-cost implementations of sub-THz wireless systems [analysisHE11, Foam-cladded, polymerblend]. RoF systems can be implemented in different ways. Herein, we consider a system with a cascaded structure, where multiple radio units (RUs) and booster units (BUs) are interconnected via low-cost PMFs and finally connected to a central unit (CU). Each RU is equipped with antennas, a power amplifier (PA), and other radio frequency (RF) components; see Section III for details. The only difference between an RU and a BU is the absence of antennas in a BU. The RU capturing or transmitting the sub-THz signals acts as an access point, and all other units between the entry/exit RU and the CU only amplify and forward the signals. All signal processing is performed at the CU. Since there are no digital signal processing units in the RUs and BUs, the system is significantly less costly than a dense deployment of radio transceivers.
@@
Fig. 1 illustrates the cascade RoF system under consideration (the different variables will be introduced later). In this system, the signal propagates over a short-range sub-THz wireless channel and over multiple dispersive fibers within the RoF system. The salient feature of this system is that the propagation distance over the sub-THz wireless channel is only a small fraction of the distance between the CU and the user equipment (UE). According to measurements reported in [Frida], state-of-the-art PMFs attenuate sub-THz signals less than 5 dB per meter in the D-band (110 to 170 GHz), while the free-space path loss in the D-band is around 70 dB on the first meter and then follows an exponential law [cai6G]. Consequently, the overall propagation attenuation is significantly less in an RoF system compared to in a pure wireless sub-THz system with the same coverage.
The power amplifiers (PAs) are identified as the dominant source of power consumption. The energy efficiency (EE) of PAs is improved if a small back-off is adopted, i.e., the PAs are allowed to work in the non-linear regime [6Gwaveform]. Therefore, at each stage of the RoF system, the signals are distorted by a non-linear (PA) and a segment of dispersive PMF, with the latter being considered a linear time-invariant (LTI) system. After multiple stages, the signal will be subjected to a cumulative distortion, which we exploit to estimate the propagation distance over the RoF stripe. In this paper, each UE is assumed to transmit data to three RoF stripes that are deployed in parallel and mounted at the ceiling, as shown in Fig 2. The CU at each stripe jointly estimates the propagation distance over the stripe, as well as the time-of-arrival (ToA) with a clock bias between the UE and the stripe. Based on estimates of ToA and propagation distance from three CUs, trilateration is used to locate the UE. A detailed geometry model and problem formulation will be presented in Section IV.
@@
Contributions: The specific contributions of this paper are (i) an analysis of the effects of non-linear PAs and dispersive PMFs connected in a cascade, and an analysis of the uplink (UL) signal; (ii) a maximum-likelihood (ML) framework and a non-linear least-squares (NLS) framework tailored to the cascade-structured RoF for estimating the entry RU and the ToA with bias; (iii) a thorough evaluation on the performance of the estimators with PAs working in the linear regime via a Cramér-Rao lower bound (CRLB) analysis; (iv) numerical results (using measured PMF dispersion characteristics and different values of non-linear factors) demonstrate that the proposed estimators perform well on the propagation distance estimation even in the presence of the effects of cascaded non-linear PAs; and (v) we demonstrate the usability of the proposed system for high-resolution positioning in indoor scenarios. To the best of our knowledge, this is the first study of an RoF system with segments of dispersive PMFs and with non-linear amplifiers connected in cascade. The proposed system and analysis have the potential to open up several novel research directions and opportunities for positioning and localization systems.
The rest of this paper is organized as follows. Section II presents related works. Section III discusses the system overview and the structure of every component within an RoF stripe. Section IV formulates the problem and the corresponding signal models. Section V evaluates the ML estimator in the linear regime by the CRLB. Section VI shows the applied channel measurement data and numerical results. Section VI-C applies the proposed algorithm in an indoor positioning use case. Section VII concludes the paper.
@@ II Related work
It is worth noting that this study considers an RoF system formed with only RF components and fibers. This implementation is novel and has not been investigated before, and it should not be confused with other RoF approaches that we introduce in the following.
In contrast to the system we consider, a traditional RoF system merges RF and optical fiber technologies, which typically contains lasers, an optical modulator/demodulator, optical amplifiers, optical fibers, a photomixer, and RF antennas. The mobile signal is carried by the modulated optical carriers and conveyed to the photomixer via a fiber. The photomixer then converts the optical carriers into RF electrical signals before the signals are transmitted by the RF antennas. Reference [Millimeter_ROF_Survey] discusses traditional RoF systems in more detail. It is broadly acknowledged that traditional RoF systems present low attenuation and high cost-efficiency due to the use of low-cost fibers whose development is reported in [analysisHE11, Foam-cladded, polymerblend]. Moreover, since the signals in such RoF systems are well confined within the fiber, these systems are not restricted in bandwidth. The work in [THzRoFExperiment] experimentally demonstrates the RoF systems in the THz frequency band, which implements a fiber-wireless seamless network in an indoor scenario. The reported results demonstrate good performance in terms of carrier-to-noise ratio. However, this type of RoF system often suffers from severe non-linearity in optical components, such as lasers [laserNonlinearity], and materials[MaterialNonlinearity], which represent the main challenge for RoF communications.
It is also worth differentiating the proposed system from radio stripes (RS), which is a commonly envisioned implementation of distributed massive MIMO (D-MMIMO). D-MMIMO distributes the antennas in a large area to provide ubiquitous service for every UE. The work in [6GCFMIMO] provides a comprehensive overview of D-MMIMO in 6G. Although the D-MMIMO offers consistently high-quality service to all UEs, accurate synchronization between APs and the CU is challenging. On the other hand, RSs connect APs sequentially in one cable/fiber to the CU [OverviewRS], which achieves simplified synchronization between APs via the shared bus. The authors in [ULPositioningRS] investigate the high-accuracy positioning capability with RSs at the sub-6 GHz band. However, previous works about RS consider that every AP is equipped with full signal processing abilities, which are considerably expensive at THz frequencies. In contrast, the unique structure considered in this paper allows dense and flexible deployment because the CU allocates all the resources and processes the signals. The studied system distinguishes itself from the distributed massive MIMO in the literature due to the absence of multiple antenna signal processing.
@@ III System Overview
Consider the system illustrated in Fig. 2, where multiple RoF stripes are deployed in parallel to serve UEs in an indoor scenario. Each RoF stripe consists of multiple RUs, BUs, and one CU at the end. One coordination unit connects the RoF stripes with the servers and coordinates different RoF stripes. All RoF stripes are installed on the ceiling at the same height. The configurations of RU, BU, CU, and PMF will be discussed below. All other components are considered to be LTI systems. In this study, we assume that those components introduce attenuation but no significant dispersion.
Figure 3: The structure of an RU.
Figure 4: Characteristics of the PA with different values of the non-linearity factor.
Figure 5: The structure of a CU.
@@ III-A Polymer microwave fiber
For a unit-length fiber segment, the channel impulse response with $L$ taps and amplitudes ($\beta_{0},\beta_{1},\cdots,\beta_{L-1}]^{\mathrm{T}}$) [^1] This model is valid in a non-linear system if the signals are properly oversampled [Thomas_ITC_2022]. We consider a third-order polynomial non-linear distortion; hence, the signals must be oversampled at least three times. can be expressed as
@@
where $n$ is the time index. Then, a fiber with $r$ unit length is equivalent to connecting $r$ LTI systems with the same impulse response $h_{n}$ in a cascade, whose impulse response is the $r$ times convolution of $h_{n}$ and is expressed as
@@
where all factors in (2) are modeled as in (1). In the frequency domain, we consider the frequency response for a set of $K$ discrete frequencies, $\boldsymbol{f}=[f_{0},\cdots,f_{K-1}]^{\mathrm{T}}$. Thus, the frequency domain representation of (2) is given by the multiplication of frequency responses. Let $\boldsymbol{H}$$H_{0},\cdots,H_{K-1}]^{\mathrm{T}}$ represent the frequency response of a unit-length fiber. Then, the frequency response of a unit-length fiber at the $k$th frequency $H_{k}$ is the Fourier transform of its impulse response given by
@@
III-B Radio Unit & Booster Unit
RUs are responsible for capturing incoming sub-THz signals, amplifying them, and forwarding them to the next component in the same stripe. Each RU consists of PMF couplers, PAs, antennas, and a switch as illustrated in Fig. 3. For the operation of the RoF system, we consider that each RU is equipped with a switch that selectively receives signals, either from the PMF or the sub-THz antennas. When the RU detects signals from the antennas, the switch disconnects the PMF input, allowing only the signals captured by the antennas to pass through. Therefore, each RU only admits signals either from the antennas or from the PMF. If there are two RUs in a stripe capturing signals, only the RU that is closer to the CU would act as an access point.
In case the RU structure does not implement a switch, thus allowing to receive signals coming from both sides, antennas and the PMF, a superposition of the received signals from different RUs would be received at the CU. This case is not addressed in the present work, as its treatment requires an entirely separate analysis.
BUs are deployed between any two adjacent RUs in an RoF stripe, as shown in Fig. 2. As previously mentioned, the BUs have the same structure as the RUs without the presence of antennas. The BUs amplify the signals and forward them to the next component.
To understand the impact of PAs, we examine both regimes of operation of PAs, linear and non-linear. The effects of non-linear PAs are characterized with a memoryless polynomial model. We use a third-order memoryless polynomial model with a factor $\lambda$, as higher-order terms are significantly less than the third-order term [nonlinearPA, majidi2014analysis]. A real-valued non-linear factor indicates only amplitude distortion, while a complex-valued non-linear factor indicates both phase and amplitude distortion. In practice, for a specific PA, $\lambda$ can be measured in a laboratory. The relationship between the input $x_{n}$ and output $y_{n}$ of a PA is given by
@@ III-B Radio Unit & Booster Unit
where $G$ is the amplification coefficient. Fig. 4 illustrates the characteristics of PA with different values of non-linear factors, where we can observe that the non-linearity arises with a signal magnitude of $\sim 0.4$ volts.
Due to the absence of a digital signal processing unit at a RU/BU, every RU/BU is incapable of processing signals. This design lowers the price of each RU/BU, but also disables the RU/BU to mitigate signal impairments introduced by PAs and PMFs. As a consequence, this results in the accumulation of signal distortion along the propagation in an RoF system.
@@ III-C Central unit
The CU is in charge of allocating resources, synchronizing the units in a stripe, and processing data. Each CU comprises PAs, mixers, couplers, ADC/DAC, phase-locked loops (PLL), and a digital signal processing unit (see Fig. 5). The CU is capable to operate in full-duplex mode to communicate with one UE simultaneously.
@@ IV Problem formulation and signal model
To localize a UE in a two-dimensional space, at a minimum, three observations are required. With the system structure of the cascaded-structured RoF, three stripes are required to receive the signals from a UE simultaneously. As shown in Fig. 6, each stripe jointly estimates the propagation distance (denoted as $r$) in the stripe as well as the ToA (denoted as $\tau$) of the incoming signals. The estimated propagation distances are used to locate three RUs that serve the UE, and the estimated ToAs are used for trilateration.
Before deriving the received signal model, we emphasize the key assumptions of the paper. The RoF stripes are assumed to be synchronized in time and phase via wired connections. From the perspective of practical operation, it is challenging to synchronize the UE and RoF stripes. Therefore, it is assumed that the UE has a clock bias and phase offset relative to all RoF stripes, which are unknown to the CU. The sub-THz signals propagate both over the air and over the fiber. Due to the sparse nature of the sub-THz wireless channel, the line-of-sight (LoS) components are considered to be dominant, whose path loss and ToA are not available to the CU. On the other hand, a fixed-length fiber segment is an LTI system whose impulse response can be measured in a laboratory. According to (1) to (4), the channel information of a fiber of any length is available to the CU. Similarly, the PA characteristics ($G,\lambda$, and noise variance at room temperature $\sigma^{2}$) are stable in the room temperature and can be measured after manufacture. In practice, a small deviation of PA characteristics does not have a significant effect, which is discussed in Section IV-D. In conclusion, the unknown parameters are i) $A$, a complex value absorbing the effects of the wireless channel and the phase offset between the UE and the RoF; ii) $\tau$, the propagation delay including ToA and a clock bias between the UE and the stripe; iii) $r$, the propagation distance in a stripe; iv) the coordinates of the UE. In particular, each CU/stripe estimates one propagation delay and propagation distance in the stripe, which are then used for trilateration to obtain the coordinates of the UE. All other variables in the signal model, as shown in the sections below, are known.
Figure 6: Positioning of the UE with three RoF stripes.
The RUs and BUs are uniformly distributed, which means that the distance between two consecutive units on an RoF stripe is fixed, and the distance between two adjacent RoF stripes is also fixed. Note that the stripes are shifted slightly relative to one another to ensure that the UE is not served by three RUs that lie on a line. This ensures that the geometric dilution of precision (GDOP) is favorable (see calculation in Section VI-C). The coordinate of the $i$th position of UE is given by $(U_{x}^{(i)},U_{y}^{(i)},U_{z})$, and the location of the CU in the $m$th stripe is given by $(P_{x}^{(m)},P_{y}^{(m)},0)$. Therefore, the distance between the UE and the RU in service at the $m$th stripe is
$d_{m,i}=\sqrt{\lvert U_{x}^{(i)}-(P_{x}^{(m)}-r)\rvert^{2}+\lvert U_{y}^{(i)}-P_{y}^{(m)}\rvert^{2}+U_{z}^{2}}.$ (5)
The distance between the UE and the ceiling, $U_{z}$, is assumed to be fixed and known to the CU. The proposed positioning algorithm is based on ToA with bias, whose GDOP of a certain UE position $(U_{x}^{(i)},U_{y}^{(i)})$ can be calculated according to the standard formulas [Shin_GDOP_2002]. The GDOP is defined as
$GDOP(i)=\sqrt{\mathrm{tr}\left(\boldsymbol{H}^{\mathrm{T}}_{\text{ToA}}(i)\boldsymbol{H}_{\text{ToA}}(i)\right)^{-1}},$ (6)
where $\boldsymbol{H}_{\text{ToA}}(i)$ is obtained from (5) and can be written as
$\boldsymbol{H}_{\text{ToA}}(i)=\begin{bmatrix}\frac{U_{x}^{(i)}-P_{x}^{(1)}+r}{d_{1,i}},\frac{U_{y}^{(i)}-P_{y}^{(1)}+r}{d_{1,i}}\\ \frac{U_{x}^{(i)}-P_{x}^{(2)}+r}{d_{2,i}},\frac{U_{y}^{(i)}-P_{y}^{(2)}+r}{d_{2,i}}\\ \frac{U_{x}^{(i)}-P_{x}^{(3)}+r}{d_{3,i}},\frac{U_{y}^{(i)}-P_{y}^{(3)}+r}{d_{3,i}}\end{bmatrix}$. (7)
It should be noted that this paper emerges as an initial study of new fundamental problems that appear when LTI systems and non-linear systems are connected in cascade. Depending on how the system is implemented and in what way the switches are configurable, $r$ might be known as a priori, in which case its estimation would be redundant. The initial working assumption of the architecture, however, was that there may not be a centrally controllable switch, such that the CU relies on the cumulative signal distortion of the cascaded LTI/non-linear propagation structure to estimate $r$.